OpenStudy (jaylelile):

Please help, I am unsure how to do this. Problem below:

OpenStudy (jaylelile):

\[V=\frac{ 1 }{ 3 } \pi r ^{2}h\]

OpenStudy (jaylelile):

solve for h

OpenStudy (pawanyadav):

You have to remain with h in RHS So first multiply by 3 both side

OpenStudy (pawanyadav):

You will get 3V=πr^2h

OpenStudy (waleed_imtiaz):

You have to separate the variable you are going to calculate, its h. so taking everything to the other side of equality will yield h. you got my point?

OpenStudy (pawanyadav):

Now you have to eliminate πr^2 from RHS. So divide by πr^2 both side.

OpenStudy (pawanyadav):

And you will get ........?

OpenStudy (jaylelile):

ohhh I understand, I was just trying to subtract everything other than the h.. so I have... \[\frac{ 3V }{ \pi r ^{2} } = h\]

OpenStudy (pawanyadav):

Correct..

OpenStudy (jaylelile):

thank you so much!

OpenStudy (pawanyadav):

You are welcome